OpenStudy (anonymous):

The width of a triangle is fixed at 24 cm. What lengths will make the perimeter greater than 92 cm?

OpenStudy (saifoo.khan):

Widths: 24 + 24 Remaining = 92 - 48 = 44

OpenStudy (saifoo.khan):

now 44 + 1 44 + 2 44 + anything +ve will make the perimeter greater than that.